# Yong Ji1, Junye Li1 and Rui Yang∗2,3

### v3

@@ 2.3 Specification property
A t.d.s. $(X,f)$ has the specification property if for any $\epsilon>0$, there is a positive integer $m=m(\epsilon)$ such that for any finite collection of intervals $I_{j}=[a_{j},b_{j}]\subset\mathbb{N}$ with $a_{j+1}-b_{j}\geq m(\epsilon)$, $j=1,\dots,k-1$, and any collection of $k$ points $\{x_{1},\dots,x_{k}\}$ of $X$, there exists a point $x\in X$ such that
@@
For instance, every topological mixing locally maximal hyperbolic set, and the full shift on $[0,1]^{\mathbb{Z}}$ have specification property. More examples of infinite entropy systems with specification property can be found in [BR23, Section IV].
A point $x\in X$ is called almost periodic if for any open neighborhood $U$ of $x$, there exists $N\in\mathbb{N}$ such that for every $n\in\mathbb{N}$, there exists an integer $k$ with $n\leq k\leq n+N$ for which $f^{k}(x)\in U$. Let $AP(X)$ denote the set of all almost periodic points of $X$. The following statement, proved in [HTZ23, Proposition 2.11] for systems with $g$-almost product property, gives a characterization of the structure of $AP(X)$.
@@
If the dynamical system $(X,f)$ has specification property, then the almost periodic set AP(X) is dense in $X$.
@@ 3 Proofs of main results
This section proves the main results of this paper. The challenging aspect lies in the estimation of the lower bound of $\overline{\rm mdim}_{M}^{S}(f,G_{K}^{\omega,U},d)$, where $S\in\{B,P,UC\}$. We will construct a Moran fractal set, which is geometrically constructed and is contained in $G_{K}^{\omega,U}$. Consequently, this set is well-suited for computing the lower bound of the metric mean dimensions of saturated sets.
@@ 3.1 Packing metric mean dimension of saturated sets
We first prove Theorem 1.1. To arrive at this goal, we divide its proof into two parts.
@@ 3.1.1 Upper bound of $\overline{\rm mdim}^{P}_{M}(f,G_{K},d)$
We employ a general method presented in [PS07, Theorem 4.1], [ZCC12, Theorem 1.1, $(2)$] and [JCZ22, Proposition 4] to derive the upper bound of $\overline{\rm mdim}_{M}^{P}(f,G_{K},d)$.
@@
[Y23, Lemma 3.1] Let $(X,f)$ be a t.d.s. and $\epsilon>0$. Let $\{F_{n}\}$ be a sequence of non-empty subsets of $X$, and let $E_{n}$ be an $(n,\epsilon)$-separated subset of $F_{n}$ with the largest cardinality. We define
@@
This essentially adheres to the proof of the upper bound of the classical variational principle (cf.[Wal82, Theorem 8.6]). ∎
Recall $G_{K}=\{x\in X:V_{f}(x)=K\}$. For $K\subset M_{f}(X)$, we set
$G^{K}:=\{x\in X:V_{f}(x)\subset K\}.$
Obviously, we have $G_{K}\subset G^{K}$.
@@
Let $(X,f)$ be a t.d.s..
$(1)$ Let $K\subset M_{f}(X)$ be a nonempty closed convex subset. Then for any $\epsilon>0$,
$h_{top}^{P}(f,G_{K},\epsilon)\leq h_{top}^{P}(f,G^{K},\epsilon)\leq\sup_{\mu\in K}\inf_{\mathrm{diam}(\xi)<\epsilon}h_{\mu}(f,\xi).$
$(2)$ If $\mu\in M_{f}(X)$, then for any $\epsilon>0$,
@@
It suffices to show $(1)$ since one can take $K=\{\mu\}$ as in $(1)$ and use $G_{\mu}=G^{\{\mu\}}$ to get the desired $(2)$.
@@
$\mathcal{R}_{K}(\delta,k):$ =$=\left\{x\in X:\ \mathcal{E}_{n}(x)\in B(K,\delta),~{}\forall n\geq k\right\},$
$\mathcal{Q}_{K}(\delta,n):$ =$=\left\{x\in X:\ \mathcal{E}_{n}(x)\in B(K,\delta)\right\},$
and
@@
Fix $x\in G^{K}$. Since we have $V_{f}(x)\subset B(K,\delta)$ for every $\delta>0$, then the index $n$ with $\mathcal{E}_{n}(x)\in M_{f}(X)\backslash B(K,\delta)$ must be finite. This shows that $x\in\mathcal{R}_{K}(\delta,k)$ for some $k$. Then, by $G^{K}\subset\cup_{k=1}^{\infty}\mathcal{R}_{K}(\delta,k)$ for every $\delta>0$, it follows from Proposition 2.1 that
@@
For any $k\geq 1$, by definitions, it holds that $\mathcal{R}_{K}(\delta_{0},k)\subset\mathcal{Q}_{K}(\delta_{0},n)$ for any $n\geq k$. Thus, we have
@@
Combining (3.1) and letting $\tau\to 0$, we conclude that $h_{top}^{P}(f,G^{K},\epsilon)\leq\Theta(K,\epsilon)$.
Next we shall prove for any $\gamma>0$, there is $\mu\in K$ such that
$\Theta(K,\epsilon)-\gamma\leq\inf_{{\rm diam}(\xi)<\epsilon}h_{\mu}(f,\xi).$
Take a strictly decreasing sequence $\{\delta_{k}\}$ of positive real numbers with $\delta_{k}\to 0$ and a strictly increasing sequence $\{n_{k}\}$ with $n_{k}\to\infty$ such that
@@
Let $F_{k}$ be a maximal $(n_{k},\epsilon)$-separated subset of $\mathcal{R}_{K}(\delta_{k},n_{k})$. We set
@@
Then $\mu_{k}\in B(K,\delta_{k})$ since $K$ is convex. In the weak-topology, choose a subsequence $k_{j}$ such that $\mu_{k_{j}}\to\mu$ as $j\to\infty$. Then $\mu\in K$ since $K$ is a closed. By Lemma 3.1, we have
$\Theta(K,\epsilon)-\gamma\leq\limsup_{j\to\infty}\frac{1}{n_{k_{j}}}\log s_{n_{k}}(\mathcal{Q}_{K}(\delta_{k},n_{k}),\epsilon)\leq\inf_{{\rm diam}(\xi)<\epsilon}h_{\mu}(f,\xi).$
@@
Thus, by the Lemma 3.2 the upper bound of the packing metric mean dimension of $G_{K}$ is given by
@@ 3.1.2 Lower bound of $\overline{\rm mdim}^{P}_{M}(f,G_{K}^{\omega,U},d)$
For any small enough $\epsilon^{\ast}>0$, we prove that
@@
Let $\{\delta_{(k,j)}\}$ be a strictly decreasing sequence indexed by $\mathcal{A}$ with $\delta_{(0,0)}<\zeta$, $\delta_{(k,j)}\to 0$ as $(k,j)\to\infty$. We shall simply write $\delta_{k}=\delta_{(k,Q_{k})}$. For $(k,j)\in\mathcal{A}$, let $\epsilon_{(k,j)}:=\frac{\epsilon^{\ast}}{2^{\iota(k,j)+2}}$, and $m_{(k,j)}:=m(\epsilon_{(k,j)})$, where $m$ is chosen as in the definition of specification property.
For any nonempty open set $U$, there exist $\tilde{\epsilon}>0$ and a point $a_{0}\in X$ such that $\overline{B(a_{0},\tilde{\epsilon})}\subseteq U$. One may assume that $\epsilon^{\ast}<\tilde{\epsilon}$. Set
@@
$(1)$ $v_{k}=\sum_{i=1}^{j_{k}}\lambda_{i}^{k}\nu_{i}^{k}$, where $\{\lambda_{i}^{k}:1\leq i\leq j_{k}\}$ are rational positive numbers, $\forall 1\leq i\leq j_{k}$ with $\sum_{i=1}^{j_{k}}\lambda_{i}^{k}=1$ and $\nu_{i}^{k}\in M_{f}^{e}(X)$;
@@
#$\#E_{k,i}\geq\exp\left[\lambda_{i}^{k}n_{k}^{\prime}\cdot\left(\underline{h}_{\nu_{i}^{k}}^{K}(f,5\epsilon^{\ast},\zeta)-\zeta\right)\right]\geq\exp\left[\lambda_{i}^{k}n_{k}^{\prime}\cdot\left(\inf_{\xi\succ\mathcal{U}}h_{\nu_{i}^{k}}(f,\xi)-\zeta\right)\right],$
@@
By the specification property, for $k\geq 1$ and each $\underline{x}_{k}=(x_{1},x_{2},\ldots,x_{j_{k}})\in E_{k,1}\times E_{k,2}\times\cdots\times E_{k,j_{k}}$, there exists $y=y(\underline{x}_{k})\in X$ so that
[…]
where $T_{1}=0$ and $T_{i}=\lambda_{1}^{k}n_{k}^{\prime}+\cdots+\lambda_{i-1}^{k}n_{k}^{\prime}+(i-1)m_{k}$ for $2\leq i\leq j_{k}$. One may assume that $d(x,y)<\frac{\epsilon^{\ast}}{2^{k+2}}$ implies $D(\delta_{x},\delta_{y})<\frac{1}{k}$ since the map $x\mapsto\delta_{x}$ is uniformly continuous.
For any two distinct $\underline{x}_{k}=(x_{1},x_{2},\ldots,x_{j_{k}})$, $\underline{x}_{k}^{\prime}=(x_{1}^{\prime},x_{2}^{\prime},\ldots,x_{j_{k}}^{\prime})$ in $E_{k,1}\times\cdots\times E_{k,j_{k}}$, one has $y(\underline{x}_{k})\neq y^{\prime}(\underline{x}_{k}^{\prime})$. Indeed, if $x_{i}\neq x_{i}^{\prime}$ for some $1\leq i\leq j_{k}$, then
@@
which implies $d_{n_{k}}(y,y^{\prime})>4\epsilon^{\ast}$. Let
@@
#$\#\mathcal{C}_{(k,Q_{k})}$ $=\Pi_{i=1}^{j_{k}}\#E_{k,i}\geq\Pi_{i=1}^{j_{k}}\exp\left[\lambda_{i}^{k}n_{k}^{\prime}\left(\inf_{\xi\succ\mathcal{U}}h_{\nu_{i}^{k}}(f,\xi)-\zeta\right)\right]$
@@
which shows that $D(\mathcal{E}_{n_{k}}(y),\sum_{i=1}^{j_{k}}\lambda_{i}^{k}\mathcal{E}_{\lambda_{i}^{k}n_{k}^{\prime}}(f^{T_{i}}(y)))<\frac{2}{k}$. Since
@@
Until now, by virtue of the specification property, we have obtained two sets $\mathcal{C}_{(k,0)}$ and $\mathcal{C}_{(k,Q_{k})}$ for each $k$. We proceed to consider the cases for $1\leq j<Q_{k}$.
Since the system has specification property, the generic point $G_{\mu}\neq\emptyset$ holds for any $\mu\in M_{f}(X)$ [PS07]. Applying this fact for any $k\geq 1$ and $1\leq j<Q_{k}$, then there exist $x_{k,j}\in G_{\mu_{(k,j)}}$ and $n_{(k,j)}\in\mathbb{N}$ such that
@@
With these data $\{\mathcal{C}_{(k,j)}\}$ and $\{n_{(k,j)}\}$, we are ready for the construction:
@@
For $(k,j)\in\mathcal{A}$ with $j=0$, we set $\mathcal{S}_{(k,j)}=\mathcal{C}_{(k,j)}$ and $c_{k,j}=n_{(k,j)}$. We divide the following two cases to define the rest $\mathcal{S}_{(k,j)}$ indexed by $j\neq 0$.
Case $1$. $j\neq Q_{k}$.
Recall that $\mathcal{C}_{(k,j)}=\{x_{k,j}\}$ is defined above. By specification property again, there exists $y_{k,j}\in X$ so that the pieces of orbits $\{x_{k,j},\cdots,f^{n_{(k,j)-1}}(x_{(k,j)})\}$ repeating $N_{(k,j)}$ times, can be $\epsilon_{(k,j)}$-traced by $y_{(k,j)}$ with the gap $m_{(k,j)}$. Now we set $\mathcal{S}_{(k,j)}=\{y_{k,j}\}$, and $c_{k,j}=n_{(k,j)}N_{(k,j)}+(N_{(k,j)}-1)m_{(k,j)}$.
@@
For any $(i_{1},\ldots,i_{N_{(k,Q_{k})}})\in\{1,2,\ldots,\#\mathcal{C}_{(k,Q_{k})}\}^{N_{(k,Q_{k})}}$, let $y(i_{1},\ldots,i_{N_{(k,Q_{k})}})\in X$ be given by the specification property, such that the orbit $\epsilon_{(k,Q_{k})}$-shadows, with the gap $m_{(k,Q_{k})}$, the pieces of orbits $\{y_{i_{l}}^{k},f(y_{i_{l}}^{k}),\ldots,f^{n_{k}-1}(y_{i_{l}}^{k})\}$, $l=1,2,\ldots,N_{(k,Q_{k})}$. We write
@@
and $c_{k,Q_{k}}=n_{k}\cdot N_{(k,Q_{k})}+(N_{(k,Q_{k})}-1)m_{(k,Q_{k})}$. Since $\mathcal{C}_{(k,Q_{k})}$ is an $(n_{k},4\epsilon^{\ast})$-separated set, this implies that $\mathcal{S}_{(k,Q_{k})}$ is a $(c_{k,Q_{k}},3\epsilon^{\ast})$-separated set. Then, by (3.1.2) and (3.9), one has the estimation for the cardinality of $\mathcal{S}_{(k,Q_{k})}$:
#$\#\mathcal{S}_{(k,Q_{k})}=\#\mathcal{C}_{(k,Q_{k})}^{N_{(k,Q_{k})}}\geq\exp\left[\sum_{0\leq j\leq Q_{k}}[(n_{(k,j)}+m_{(k,j)})N_{(k,j)}]\cdot(1-\zeta)^{2}\cdot(H-3\zeta)\right].$ (3.10)
Similar to the arguments in (3.8) one has
$D(\mathcal{E}_{c_{k,j}}(y_{k,j}),\mu_{(k,j)})<\frac{2}{k}+\delta_{(k,j)},\ \forall\ y_{k,j}\in\mathcal{S}_{(k,j)}~{}\text{with}~{}k\geq 1\ \text{and }1\leq j<Q_{k},$
@@
$D(\mathcal{E}_{c_{k,j}}(y),\mu_{(k,j)})<\frac{5}{k}+2\delta_{(k,j)},\ \forall\ y\in\mathcal{S}_{(k,j)}\ \text{with }k\geq 1,\ 1\leq j\leq Q_{k}.$ (3.11)
@@
One may regard this process as an analogue of the construction of the middle three Cantor set of the unit interval $[0,1]$.
@@
[…]
Note that $M_{(k,j)}^{N_{(k,j)}}=M_{(k,j)}$ and the number of $\mathcal{S}_{(k,Q_{k})}$ given in formula (3.10) has the estimation:
#$\#\mathcal{S}_{(k,Q_{k})}\geq\exp\left[(M_{(k,Q_{k})}-M_{(k-1,Q_{k-1})})(1-\zeta)^{2}(H-3\zeta)\right].$ (3.12)

### v4

@@ 2.3 Specification property of TDSs
One says that a TDS $(X,d,f)$ has the specification property if for any $\epsilon>0$, there is a positive integer $m=m(\epsilon)$ such that for any finite collection of intervals $I_{j}=[a_{j},b_{j}]\subset\mathbb{N}$ with $a_{j+1}-b_{j}\geq m(\epsilon)$, $j=1,\dots,k-1$, and any collection of $k$ points $\{x_{1},\dots,x_{k}\}$ of $X$, there exists a point $x\in X$ such that
@@
For instance, every topological mixing locally maximal hyperbolic set, and the full shift on $0,1]^{\mathbb{Z}}$ have the specification property. More examples of infinite entropy systems with the specification property can be found in [BR23, Section IV].
A point $x\in X$ is called almost periodic if for any open neighborhood $U$ of $x$, there exists $N\in\mathbb{N}$ such that for every $n\in\mathbb{N}$, there exists an integer $k$ with $n\leq k\leq n+N$ for which $f^{k}(x)\in U$. Denote by $AP(X)$ the set of all almost periodic points of $X$.
@@
[HTZ23, Proposition 2.11] If a TDS $(X,d,f)$ has the specification property, then the almost periodic set AP(X) is dense in $X$.
@@ 3 Proofs of main results
In this section, we prove the main results of this paper.
@@ 3.1 Packing metric mean dimension of saturated sets
To prove Theorem 1.1, we divide its proof into two parts.
@@ 3.1.1 Upper bound of $\overline{\rm mdim}^{P}_{M}(f,G_{K},d)$
[Y23, Lemma 3.1] Let $(X,d,f)$ be a TDS and $\epsilon>0$. Let $\{F_{n}\}$ be a sequence of non-empty subsets of $X$, and let $E_{n}$ be an $(n,\epsilon)$-separated subset of $F_{n}$ with the largest cardinality. We define
@@
It essentially follows from the proof of the upper bound for the classical variational principle (cf.[Wal82, Theorem 8.6]). ∎
@@
Let $(X,d,f)$ be a TDS, and let $K\subset M_{f}(X)$ be a non-empty closed convex subset. Then for any $\epsilon>0$,
$h_{top}^{P}(f,G_{K},\epsilon)\leq h_{top}^{P}(f,G^{K},\epsilon)\leq\sup_{\mu\in K}\inf_{\mathrm{diam}(\xi)<\epsilon}h_{\mu}(f,\xi),$
where $G^{K}:=\{x\in X:V_{f}(x)\subset K\}.$
Consequently, for every $\mu\in M_{f}(X)$ and $\epsilon>0$,
@@
It suffices to show the inequality
$h_{top}^{P}(f,G^{K},\epsilon)\leq\sup_{\mu\in K}\inf_{\mathrm{diam}(\xi)<\epsilon}h_{\mu}(f,\xi).$
@@
$\mathcal{R}_{K}(\delta,k):$ $=\left\{x\in X:\ \mathcal{E}_{n}(x)\in B(K,\delta),~\forall n\geq k\right\},$
$\mathcal{Q}_{K}(\delta,n):$ $=\left\{x\in X:\ \mathcal{E}_{n}(x)\in B(K,\delta)\right\},$
and define
@@
For every $x\in G^{K}$, we have $V_{f}(x)\subset B(K,\delta)$ for every $\delta>0$. Then the cardinality of $n$ such that $\mathcal{E}_{n}(x)\in M_{f}(X)\backslash B(K,\delta)$ must be finite. This shows that $x\in\mathcal{R}_{K}(\delta,k)$ for some $k$. By $G^{K}\subset\cup_{k=1}^{\infty}\mathcal{R}_{K}(\delta,k)$ for every $\delta>0$, it follows from Proposition 2.1 that
@@
For any $k\geq 1$, it holds that $\mathcal{R}_{K}(\delta_{0},k)\subset\mathcal{Q}_{K}(\delta_{0},n)$ for any $n\geq k$. Thus, we have
@@
Using (3.1) and letting $\tau\to 0$, we conclude that $h_{top}^{P}(f,G^{K},\epsilon)\leq\Theta(K,\epsilon)$.
Take a strictly decreasing sequence $\{\delta_{k}\}$ with $\delta_{k}\to 0$ and a strictly increasing sequence $\{n_{k}\}$ with $n_{k}\to\infty$ such that
@@
Let $F_{k}$ be a $(n_{k},\epsilon)$-separated subset of $\mathcal{R}_{K}(\delta_{k},n_{k})$ with the largest cardinality. We set
@@
Then $\mu_{k}\in B(K,\delta_{k})$ since $K$ is convex. Now choose a subsequence $\{n_{k_{j}}\}$ such that $\mu_{k_{j}}\to\mu$ as $j\to\infty$. Then $\mu\in K$ since $K$ is a closed. By Lemma 3.1, we have
$\Theta(K,\epsilon)-\gamma\leq\limsup_{j\to\infty}\frac{1}{n_{k_{j}}}\log s_{n_{k_{j}}}(\mathcal{Q}_{k_{j}}(\delta_{k_{j}},n_{k_{j}}),\epsilon)\leq\inf_{{\rm diam}(\xi)<\epsilon}h_{\mu}(f,\xi).$
@@
Thus, by Lemma 3.2 the upper bound of $\overline{\rm mdim}_{M}^{P}(f,G_{K},d)$ is bounded above by
@@ 3.1.2 Lower bound of $\overline{\rm mdim}^{P}_{M}(f,G_{K}^{\omega,U},d)$
The challenging direction is how to estimate the lower bound of $\overline{\rm mdim}^{P}_{M}(f,G_{K}^{\omega,U},d)$. This shall be done by constructing a geometric Moran fractal set that is contained in $G_{K}^{\omega,U}$.
For any sufficiently small $\epsilon^{\ast}>0$, we prove that
@@
Let $\{\delta_{(k,j)}\}$ be a strictly decreasing sequence indexed by $\mathcal{A}$ with $\delta_{(0,0)}<\zeta$, $\delta_{(k,j)}\to 0$ as $(k,j)\to\infty$. We shall simply write $\delta_{k}=\delta_{(k,Q_{k})}$. For $(k,j)\in\mathcal{A}$, let $\epsilon_{(k,j)}:=\frac{\epsilon^{\ast}}{2^{\iota(k,j)+2}}$, and $m_{(k,j)}:=m(\epsilon_{(k,j)})$ be chosen as in the definition of the specification property.
For any non-empty open set $U$, there exist $0<\tilde{\epsilon}<\epsilon^{*}$ and a point $a_{0}\in X$ such that ${\overline{B}(a_{0},\tilde{\epsilon})}\subseteq U$. Set
@@
$(1)$ $v_{k}=\sum_{i=1}^{j_{k}}\lambda_{i}^{k}\nu_{i}^{k}$, where $\nu_{i}^{k}\in M_{f}^{e}(X)$, $\lambda_{i}^{k}\in\mathbb{Q}^{+}$ for every $1\leq i\leq j_{k}$, and $\sum_{i=1}^{j_{k}}\lambda_{i}^{k}=1$;
@@
$E_{k,i}\geq\exp\left[\lambda_{i}^{k}n_{k}^{\prime}\cdot\left(\underline{h}_{\nu_{i}^{k}}^{K}(f,5\epsilon^{\ast},\zeta)-\zeta\right)\right]\geq\exp\left[\lambda_{i}^{k}n_{k}^{\prime}\cdot\left(\inf_{\xi\succ\mathcal{U}}h_{\nu_{i}^{k}}(f,\xi)-\zeta\right)\right],$
@@
By the specification property, for $k\geq 1$ and each $\underline{x}_{k}=(x_{1},x_{2},\ldots,x_{j_{k}})\in E_{k,1}\times E_{k,2}\times\cdots\times E_{k,j_{k}}$, there exists $y=y(\underline{x}_{k})\in X$ such that
$d\left(f^{T_{i}+t_{i}}(y),f^{t_{i}}(x_{i})\right)<\epsilon_{(k,Q_{k})}<\frac{\epsilon^{\ast}}{2^{k+2}},\ 1\leq i\leq j_{k},\ 0\leq t_{i}\leq\lambda_{i}^{k}n_{k}^{\prime}-1,$
where $T_{1}=0$ and $T_{i}=\lambda_{1}^{k}n_{k}^{\prime}+\cdots+\lambda_{i-1}^{k}n_{k}^{\prime}+(i-1)m_{k}$ for $2\leq i\leq j_{k}$. For any two distinct $\underline{x}_{k}=(x_{1},x_{2},\ldots,x_{j_{k}})$, $\underline{x}_{k}^{\prime}=(x_{1}^{\prime},x_{2}^{\prime},\ldots,x_{j_{k}}^{\prime})$ in $E_{k,1}\times\cdots\times E_{k,j_{k}}$, one has $y(\underline{x}_{k})\neq y^{\prime}(\underline{x}_{k}^{\prime})$. Indeed, if $x_{i}\neq x_{i}^{\prime}$ for some $1\leq i\leq j_{k}$, then
@@
which implies $d_{n_{k}}(y,y^{\prime})>4\epsilon^{\ast}$.
Let
@@
$\#\mathcal{C}_{(k,Q_{k})}$ $=\Pi_{i=1}^{j_{k}}\#E_{k,i}\geq\Pi_{i=1}^{j_{k}}\exp\left[\lambda_{i}^{k}n_{k}^{\prime}\left(\inf_{\xi\succ\mathcal{U}}h_{\nu_{i}^{k}}(f,\xi)-\zeta\right)\right]$
@@
which shows that $D(\mathcal{E}_{n_{k}}(y),\sum_{i=1}^{j_{k}}\lambda_{i}^{k}\mathcal{E}_{\lambda_{i}^{k}n_{k}^{\prime}}f^{T_{i}}(y))<\frac{2}{k}$. Since the map $x\mapsto\delta_{x}$ is uniformly continuous, we may assume that for any $x,y\in X$ with $d(x,y)<\frac{\epsilon^{\ast}}{2^{k+2}}$ implies $D(\delta_{x},\delta_{y})<\frac{1}{k}$. Since
@@
Until now, by virtue of the specification property, we have obtained two sets $\mathcal{C}_{(k,0)}$ and $\mathcal{C}_{(k,Q_{k})}$ for each $k\geq 1$. We proceed to consider the cases for $1\leq j<Q_{k}$.
Since the system has the specification property, the generic point $G_{\mu}\neq\emptyset$ holds for all $\mu\in M_{f}(X)$ [PS07]. Using this fact for any $k\geq 1$ and $1\leq j<Q_{k}$, there exist $x_{k,j}\in G_{\mu_{(k,j)}}$ and $n_{(k,j)}\in\mathbb{N}$ such that
@@
With these data $\{\mathcal{C}_{(k,j)}\}$ and $\{n_{(k,j)}\}$, we are ready to proceed the next construction:
@@
For $(k,0)\in\mathcal{A}$, we set $\mathcal{S}_{(k,0)}:=\mathcal{C}_{(k,0)}$ and $c_{k,0}:=n_{(k,0)}$. We divide the following two cases to define the rest $\mathcal{S}_{(k,j)}$ indexed by $j\neq 0$.
Case $1$. $1\leq j<Q_{k}$.
Recall $\mathcal{C}_{(k,j)}=\{x_{k,j}\}$. By the specification property again, there exists $y_{k,j}\in X$ such that the pieces of orbits $\{x_{k,j},\cdots,f^{n_{(k,j)-1}}(x_{(k,j)})\}$ repeating $N_{(k,j)}$ times, can be $\epsilon_{(k,j)}$-traced by $y_{(k,j)}$ with the gap $m_{(k,j)}$. Now we set $\mathcal{S}_{(k,j)}=\{y_{k,j}\}$, and $c_{k,j}=n_{(k,j)}N_{(k,j)}+(N_{(k,j)}-1)m_{(k,j)}$.
@@
For any $(i_{1},\ldots,i_{N_{(k,Q_{k})}})\in\{1,2,\ldots,\#\mathcal{C}_{(k,Q_{k})}\}^{N_{(k,Q_{k})}}$, let $y(i_{1},\ldots,i_{N_{(k,Q_{k})}})\in X$ be a point given by the specification property, such that the orbit $\epsilon_{(k,Q_{k})}$-shadows, with the gap $m_{(k,Q_{k})}$, the pieces of orbits $\{y_{i_{l}}^{k},f(y_{i_{l}}^{k}),\ldots,f^{n_{k}-1}(y_{i_{l}}^{k})\}$, $l=1,2,\ldots,N_{(k,Q_{k})}$. We write
@@
and $c_{k,Q_{k}}=n_{k}\cdot N_{(k,Q_{k})}+(N_{(k,Q_{k})}-1)m_{(k,Q_{k})}$. Since $\mathcal{C}_{(k,Q_{k})}$ is an $(n_{k},4\epsilon^{\ast})$-separated set, this implies that $\mathcal{S}_{(k,Q_{k})}$ is a $(c_{k,Q_{k}},3\epsilon^{\ast})$-separated set. Then, by (3.1.2) and (3.9), the cardinality of $\mathcal{S}_{(k,Q_{k})}$ is bounded blew by
$\mathcal{S}_{(k,Q_{k})}=\#\mathcal{C}_{(k,Q_{k})}^{N_{(k,Q_{k})}}\geq\exp\left[\sum_{0\leq j\leq Q_{k}}[(n_{(k,j)}+m_{(k,j)})N_{(k,j)}]\cdot(1-\zeta)^{2}\cdot(H-3\zeta)\right].$ (3.10)
Similar to the arguments in (3.8), one has
[…]
@@
[…]
@@
This step can be regarded as an analogue of the construction of the middle-third Cantor set of $[0,1]$.
@@
$M_{(k,j)}^{i}$ $=M_{(k,j)^{\ast}}+i(m_{(k,j)}+n_{(k,j)}),\ 0\leq i\leq N_{(k,j)}.$
Note that $M_{(k,j)}^{N_{(k,j)}}=M_{(k,j)}$ and the number of $\mathcal{S}_{(k,Q_{k})}$ given in (3.10) has the estimation:
$\mathcal{S}_{(k,Q_{k})}\geq\exp\left[(M_{(k,Q_{k})}-M_{(k-1,Q_{k-1})})(1-\zeta)^{2}(H-3\zeta)\right].$ (3.12)